OpenStudy (anonymous):

Find the rate of change between the two points: (6, 8) and (5, 6) where x is in days and y is in pounds.

OpenStudy (anonymous):

any help.......?

OpenStudy (anonymous):

change of rate = change in pounds/change in days dy/dx = (6-8)/(5-6) dy/dx = (-2)/(-1) dy/dx = 2 pounds per day

OpenStudy (anonymous):

thanks so much ;o is that it??

OpenStudy (anonymous):

Yups. :)

OpenStudy (anonymous):

Thank you ^-^

OpenStudy (anonymous):

Anytime :)